A man is 10 years older than his son. In ten years, he will be twice as old as his son will then be. How old is each
Answer by Edwin McCravy(20056)

A man is 10 years older than his son.
What??? The man became a father at age 10?????

Surely you miscopied the problem.

In ten years, he will be twice as old as his son will then be.
How old is each.
If that were the bizarre case, then

son's age now = x
father's age now = x + 10

son's age in 10 years = x + 10
father's age in 10 years = (x + 10) + 10 or x + 10 + 10 or x + 20

So 

x + 20 = 2(x + 10)

x + 20 = 2x + 20

    -x = 0

     x = 0

son's age now = x = 0
father's age now = x + 10 = 0 + 10 = 10

So the 10-year old boy just became a father, for the baby is
0 years old (just born).  And in ten years the 10-year-old boy 
father will be 20 years old and the son will be 10.

That works, but surely the problem is botched.

Edwin